OpenStudy (anonymous):

Find the equation of the tangent line to y=ln(x-1) at the point (2,0). -Thanks

OpenStudy (anonymous):

y'(x)=1/(x-1) y'(2) = 1 So line equation is: y=x-2 If need more explain tell me,:)

OpenStudy (anonymous):

Ohh hmm, maybe the answer is wrong. It says y=x+2

OpenStudy (anonymous):

do you know why>?

OpenStudy (experimentx):

the answer must be wrong ... i guess,

OpenStudy (anonymous):

So what is the general method to do this?

OpenStudy (anonymous):

I think i am right

OpenStudy (anonymous):

Haha thanks but can you explain a little more about how you got this answer?

OpenStudy (anonymous):

Like i understand when you get y=1/x-1 But then after, what happens>?

OpenStudy (anonymous):

y'(2) gives the slope. standart equation of line: y-y0=m(x-x0) where (x0,y0) is the point of the line. rest is easy

OpenStudy (anonymous):

Okay, so after you find the derivative, to find the slope at that point, you sub in x=2?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

you welcome